an cject could actelusti Centinueln tlo.omls he eng Bwl take to Reach 2
Answer #1

a) . t= v/a = 334/10 = 33.4 sec

b) t= 11200/10 = 1120 sec = 18.67 min .  

c) 10% of speed of light = 3x107 m/s

time = 3x107/10 = 3x106 sec
= 34.7 days .